# The Role of Task Complexity in Emergent Abilities of Small Language Models

### v1

@@ Abstract
We investigate the relationship between task complexity and the minimum model size required for learning specific tasks in small transformer models. We focus on the ListOps dataset, consisting of nested math operations. We define the task complexity as the Kolmogorov complexity (KC) of the code solving the task, using a rough proxy for KC. We find a power-law relation between KC and parameters required to learn, suggesting number of parameters to learn harder task increases almost cubic in KC. On individual math operations, sum mod 10 is hardest to learn. Surprisingly, when combining tasks, we observe that sum is learned earlier and with fewer parameters when trained alongside max and median. Analyzing the model, we find strong evidence that models trained on sum alone and models trained jointly converge to different algorithms. Concretely, the sum alone model doesn’t seem to have learned number properties in the embedding layer, likely memorizing the sum table. In contrast, models trained on three tasks (maximum, median and sum) reveals that joint training results in clear number-like properties. Finally, we also find evidence that the sum-only model utilizes its feedforward layer more than the jointly trained model. Conversely, the attention layer in the joint model is activated more than the sum model. Our findings suggest there is another dimension to emergent abilities in language models, namely the algorithms being learned, potentially impacting scaling laws.
@@
1 Related works
Scaling laws for language models characterize how validation loss scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design. However, this perspective overlooks critical aspects of learning. Our results show that the composition of the training data—specifically, the mix of tasks—can significantly alter what and how a model learns. Even models with identical sizes can follow different learning trajectories depending on the training setup, suggesting that task structure shapes the internal algorithms that emerge during training.
In general, language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study. Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms. While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively. Furthermore, it allows us to tune the task difficulty by combining different operations.
Here, we analyze how language models learn arithmetic operations by training small models on the ListOps dataset. The ListOps dataset allows us to measure model accuracy explicitly by asking the model to solve mathematical equations. Using this bottom-up approach, we show that once the number of trainable parameters in the model reaches a critical point, the model is able to learn the task. When training models on datasets based on different levels of operation difficulty, this critical point shifts according to the complexity of the task. Contradictorily, we find that combining operations makes the task easier for the model than individual operations, suggesting that task diversity plays a significant role.
lubana2024percolation discuss that structure learning drives emergent abilities. Seems very relevant. Emergence of hidden capabilities park2024emergence Another grokking paper liu2022towards claims grokking is between “ comprehension” and “memorization”. We may have more evidence for that.
@@ 2 Introduction
Scaling laws for language models kaplan2020scaling characterize how performance scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design. Moreover, study of emergent abilities in LLM showed wei2022emergent that various capabilities emerge at different model sizes. Both of these directions suggest an intimate and universal relations between model size and the emergence of abilities. However, these views miss important details about the role of training curriculum.
In general language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study. Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms. While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively. Furthermore, it allows us to tune the task difficulty by combining different operations.
There is always more than one way to solve a task. Each solution might require dramatically different amount of compute and memory resources. If models learn to implement specific solution algorithms, variability in resource requirements should be reflected in some way in the model size or other metrics. Oftentimes, when humans try to solve a task, the simplest solution (meaning, easiest to find) is a brute-force one and not particularly efficient. Given more thought, knowledge and experience, we often manage to find more efficient solutions. So a key question to ask is whether the same type of algorithmic change can be observed in machine learning models and how it can be controlled.
How do we know which solution path a model took? We explore this problem through the lens of basic math operations, using the setting of the ListOps dataset (nangia2018listops). We show strong evidence for learning curriculum dramatically changing the algorithms learned by the models. Models of the same size may follow different learning paths depending on the training setup, suggesting that the tasks themselves influence the internal algorithms that emerge during training.
It is worth noting that our observations point to more than just task complexity being the driving force. Complexity of a task may be defined using Kolmogorov complexity (KC) kolmogorov1965three; li2008introduction, defined as the length of the shortest code producing a desired output. What we show is that models may require vastly different minimum sizes to learn the same task, depending on the training curriculum. Additionally, we show strong evidence of a clear distinction between the internal algorithm learned by the models using different curricula. Such a distinct algorithm could represent learning different algorithms to solve the same problem, with each algorithm having a different resource requirement, e.g. brute-force versus an efficient code. Hence we argue the model size depends more on the algorithmic complexity, or description length of a specific algorithm, rather than the KC of the task. This also raises an important question: what training regimes are most effective at inducing the learning of more efficient algorithms?
Using the ListOps dataset as our experimental framework, we investigate how small transformer models learn nested mathematical operations—specifically maximum (MAX), minimum (MIN), median (MED), and sum modulo $n$ (SUM). We train the models on a single or a mixture of these operations. We make the following observations:
@@
    A clear hierarchy in the model size vs learning transition: MAX and MIN are easiest to learn. Introducing MED and all combinations of MAX, MIN and MED require significantly more parameters to learn (2$\sim$ 2$\times$ 2 × for $n=26$). Next, adding SUM (mod $n$) to the mix makes it again significantly harder (1.5$\sim$ 1.5$\times$ 1.5 × for $n=26$).
@@
    SUM is the hardest: Paradoxically, the clear outlier as the hardest task to learn was SUM alone, not the more complex mixture of SUM and other operations.
@@
    Strong evidence for algorithmic diversity: The embedding layer of the pure SUM models, suggests models not learning the properties of numbers, whereas, mixed SUM models show strong number features including ordering and parity, with parity only observed in mixed SUM tasks.
@@
    Pretraining with MAX+MED makes SUM easier: Switching from MAX+MED to SUM has the same effect as mixed task, with much smaller models being able to learn SUM.
We hypothesize that compared to the pure SUM task, the mixed SUM training makes it easier for the models to discover number properties of the symbols. This seems to lead the mixed SUM models learning an entirely different algorithm than the pure SUM model. It may be that the pure SUM models are memorizing the table of additions. Indeed, we observe a similar difficulty to learning pure SUM for models learning a scrambled SUM dataset, where we create a scrambled sum table with the right hand side of $A+B=C$ being shuffled. In the scrambled SUM task, any number relation between the symbols has been erased and the easiest way to learn it should be to memorize the sum table, rather than overfit the random patterns in the sum table.
These results provide compelling evidence that joint training guides models towards finding alternative, more efficient solutions. They suggest there is more nuance to the scaling laws and they can be significantly affected by the training curriculum and strategies.
@@ 3 Methodology
Choice of task
ListOps consists of nested mathematical equations involving operations. Our setup uses MAX, MIN, MED (median), and SUM (sum modulo $n$) applied to numbers (0-$n-1$). We conduct experiments using both $n=10$ and $n=26$, in which case we use capital letters A…Z to represent $0\dots 25$ (Appendix D.1). Choosing ListOps was motivated by several key factors that align with our research objectives:
  1. 1.
    Procedural Generation: ListOps is a synthetically generated dataset, allowing us to create a large volume of examples with controlled difficulty.
  2. 2.
    Exact Evaluation: The mathematical nature of ListOps operations ensures unambiguous, exact evaluation of model outputs.
  3. 3.
    Adjustable Task Complexity: ListOps offers a framework to modulate problem complexity by combining different operations and adjusting nesting depths.
  4. 4.
    Inter-task Relationships: The dataset’s multiple operations provide an opportunity to explore task synergies and interference.
Dataset Description
@@
Tokenization and CoT:
We employ a character-based tokenization strategy for processing ListOps expressions. (Appendix D.2). We find that directly solving nested ListOps in one step can be quite challenging for transformer model (Fig. 22) Even with a maximum of three nesting levels with three operands (inputs) we find that GPT models with over 10 million parameters still fail to learn the task. To enhance model performance, particularly on more complex operations like sum modulo 10, we introduced a chain of thought (CoT) approach in our training data (Appendix D.3): s(%12(%34))>(%127)>0=0e, where the ‘>’ token means one step of CoT, wherein we solve the right-most, inner-most parenthesis. This is similar to the scratchpad used in lee2024teaching, among others. The tasks become increasingly challenging to learn with increasing nesting, number of operands, and length in a predictable manner (Appendix Fig. 27).
Train Test Split:
The CoT could result in many patterns appearing in one equation. To ensure that all the patterns in the test set are not also present in the training data, we select 100 triplets of numbers, e.g. $749$, to make an exclusion set. The training data is chosen from equations which never encounter the triplets in the exclusion set, and the test set is chosen such that each equation contains at least one excluded triplet.
Performance Evaluation:
We randomly select 1000 equations from the held-out test set. The model is prompted to generate solutions character-by-character, starting from the equation prompt (e.g., s(+3(-749))>). We use the output to produce two metrics: Loss: using cross-entropy computed for every character of the output. Accuracy: based on the number of correct answers evaluated using only the final answer, which we define as the first character after the first ‘$=$’ symbol.
Model Architecture
For our experiments, we employed a series of tiny GPT models, inspired by the nanoGPT architecture (karpathy2022nanogpt). Each model in our study uses a single attention head. We set the feedforward hidden dimension to four times the embedding dimension, a common practice in transformer architectures, providing sufficient complexity in the feedforward networks while keeping the model size constrained. By varying the embedding size and the number of transformer layers, we can create a range of models with different parameter counts.
@@ 4 Results
To understand how language models acquire mathematical abilities, we focus on accuracy as the primary metric for observing the emergence of learning, following the approach of the emergent abilities literature as in wei2022emergent. While some subsequent studies have questioned the concept of emergence by examining other metrics (schaeffer2024emergent), we argue that this critique overlooks similar patterns observed in physics: In a phase transition, only some quantities may exhibit discontinuous changes. A metric that can capture the emergence of the new phase is called the “order parameter”. Hence, we choose to use accuracy as our order parameter.
We run experiments with embedding dimension $n_{embed}$ between 4–128, and number of layers 1–4. In most settings, the single-layer transformers never reached 100% accuracy, so we decided to omit them. Similar to the emergence literature, we observe the total number of parameters to be the strongest indicator of accuracy, rather than $n_{embed}$ or depth. We mainly show results for modulo 26 experiments, as the patterns are more conclusive, but we also observed similar patterns for mod 10 (Appendix A.2).

### v2

# Small Models, Smarter Learning: The Power of Joint Task Training

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Scaling laws predict language model performance as a function of model size (Hoffmann et al., 2022), compute (Muennighoff et al., 2023), and dataset size (Hestness et al., 2017), but treat training data as homogeneous and do not explain how task composition shapes learning efficiency. Prior work has shown that multi-task learning improves generalization (Caruana, 1997; Ruder, 2017) and that joint arithmetic training can benefit individual tasks (Lee et al., 2024; Abedsoltan et al., 2025). However, a systematic understanding of when joint training helps versus fails—and by how much—remains lacking.
[…]
Mathematical tasks provide an ideal testbed for this investigation: they allow precise control over task difficulty, quantitative evaluation of accuracy, and systematic combination of operations (Nangia and Bowman, 2018; Nye et al., 2021; Dziri et al., 2023). We use the ListOps dataset (Nangia and Bowman, 2018), which consists of nested operations such as MAX(2,3,ADD(4,5)), extended with modular arithmetic and permutation groups.
Summary of contributions.
We study how small transformer models learn nested operations—maximum (MAX), minimum (MIN), median (MED), modular addition (ADD), modular product (PROD), and alternating addition-subtraction (NADD: $\sum_{i}(-1)^{i}x_{i}$)—both in isolation and jointly. We further extend to permutation groups, where products can be decomposed into sub-block operations. Our main findings are:
@@
    Task difficulty hierarchy: MAX/MIN are easiest; MED and PROD (non-prime modulus) are intermediate; ADD, NADD, and PROD (prime modulus) are hardest, requiring an order of magnitude more parameters.
@@
[…]
@@
    Synergies that fail: Pairing structurally similar hard tasks (ADD+NADD) provides no benefit. Nor does pairing ADD with PROD when the modulus is prime. This indicates that task synergy depends on shared computational primitives, not superficial similarity.
@@
    Mechanistic evidence: PCA of learned embeddings shows that successful joint training induces structured number representations (ordering, parity, divisibility) absent in single-task models.
  5. 5.
[…]
These results demonstrate that joint training can guide models toward more efficient solutions, revealing structure in scaling laws that depends critically on task composition. Understanding these synergies has practical implications: if certain task combinations make models “smarter” by inducing better representations, this could inform curriculum design for pretraining—a form of pre-pretraining that reduces the resources needed for downstream tasks.
2 Related Work
[…]
Curriculum learning. Presenting tasks in order of difficulty improves learning (Bengio et al., 2009; Yin et al., 2024; Kim and Lee, 2024). Our transfer experiments (Section 4.3) relate to curriculum learning, showing that pretraining on easy tasks enables learning hard tasks at smaller model sizes. However, our focus is on simultaneous joint training rather than sequential curricula.
Grokking and mechanistic interpretability. Work on grokking (Power et al., 2022; Liu et al., 2022) and mechanistic interpretability (Nanda et al., 2023) has shown that models solving modular arithmetic develop structured internal representations. We build on these insights by using embedding analysis to understand why joint training helps: successful task pairings induce structured embeddings (ordering, parity) that single-task training does not. Our setup differs from grokking studies in that we vary model size at fixed dataset size, rather than observing delayed generalization at fixed model size.
We provide the first systematic map of task compatibility for arithmetic operations, extending to permutation groups and confirming robustness to architectural choices (Appendix H).
@@ 3 Methodology
Choice of task.
ListOps consists of nested mathematical equations involving operations. Our setup uses MAX, MIN, MED (median), ADD (addition modulo $n$), PROD (product modulo $n$) and NADD (alternating addition and subtraction modulo $n$) applied to numbers (0 to n1n-1). We conduct experiments using $n=10$ to $n=226$ (Appendix F.3.1). ListOps provides procedural generation with controlled difficulty, exact evaluation, adjustable complexity via nesting depth, and systematic investigation of inter-task relationships.
Dataset description.
We use a simplified functional notation of the form $f(x,y,\ldots)$. For example, max_10(3,min_10(7,4,9))=4<eos>, where <eos> denotes the end token. The subscript in the function name indicates the modulo operation e.g., max_10 represents the max function modulo $10$.
Tokenization and CoT.
We use word-level tokenization: each operation (e.g., add_20), each number (e.g., 13), and each symbol ((, ), =, <eos>) is a single token (Appendix F.3.2). Since directly solving nested expressions proves challenging (Fig. 24), we use chain-of-thought (CoT) that decomposes expressions into intermediate steps (Appendix F.3.3): add_10(9 add_10(3 4))>add_10(9 7)>6=6<eos>, where > indicates solving the innermost operation. This follows the scratchpad approach of Lee et al. (2024). Task difficulty increases with the number of operands and nesting depth (Appendix Fig. 48).
Train-Test split.
The experiments were conducted on data generated with a maximum nesting depth of 2 and up to 3 operands per operation. Equations are constructed from digit sequences forming duplets and triplets (e.g., ”1,0”, ”0,0,1”). Since the CoT format can produce many intermediate patterns within a single equation, we implement a rigorous train-test split to prevent data leakage: we designate 100 of the 1000 possible triplets (e.g. ”749”) as an exclusion set. Training equations never contain any excluded triplet, while test equations each contain at least one excluded triplet, ensuring that test set patterns are novel.
Permutation groups.
To test generality beyond arithmetic, we extend ListOps to permutation groups. By Cayley’s theorem, every finite group is isomorphic to a permutation subgroup, making this a natural generalization. We construct $6\times 6$ block-diagonal permutation matrices (two $3\times 3$ blocks) and define three operations: OP (full matrix, 36 elements), OP_TOP and OP_BOTTOM (individual blocks, 6 elements each). See Fig. 9a and Appendix Figs. 1513 for details.
Performance evaluation.
We randomly select 1000 equations from the held-out test set. The model is prompted to generate solutions character-by-character, starting from the equation prompt (e.g., add_10(3,min_10(7,4,9))>). We use the output to produce two metrics: Loss: using cross-entropy computed for every character of the output. Accuracy: based on the number of correct answers evaluated using only the final answer, which we define as the first character after the first ‘==’ symbol.
Figure 2: Joint training with compatible tasks reduces the model size needed to learn. Each panel shows accuracy vs. model size for operations trained in isolation ($\bullet$), with an easier task ($\blacksquare$), or with a harder task ($\blacktriangle$). Left: ADD benefits from PROD ($2.5\times$ smaller models) but not from NADD. Center: NADD benefits from PROD but not from ADD. Right: PROD benefits from MAX ($2.6\times$ smaller) but not from ADD+MAX. Dashed lines show logistic fits; legends report the 80% accuracy transition point.
Model architecture.
For our experiments, we employed a series of tiny GPT models, inspired by the nanoGPT architecture (Karpathy, 2022). Unlike the standard sequential transformer, we implemented a recurrent variant in which a single transformer block was iteratively applied by feeding back its output as the next input (see the architecture Appendix Fig. 23). This recurrent design improved learning efficiency and yielded more structured embedding patterns (see results from traditional deep transformer models in the Appendix F). Each model in our study uses a single attention head. We set the feedforward hidden dimension to four times the embedding dimension, a common practice in transformer architectures, providing sufficient complexity in the feedforward networks while keeping the model size constrained. By varying the embedding size we create a range of models with different parameter counts. Ablation studies with 1–4 attention heads show no qualitative change in results (Appendix H), and experiments with standard deep transformers (2–4 layers) yield similar patterns, though with less structured embeddings (Appendix F).
Learning transition.
We define the learning transition as the minimum model size at which a task can be learned. Operationally, we fit a logistic function to accuracy vs. model size and define the transition point as the model size reaching 80% of the asymptotic accuracy. We chose 80% rather than 50% because the latter corresponds to models that have not yet reliably mastered the task; 80% ensures the model has learned a generalizable solution rather than partially memorizing patterns. The transition curves are remarkably consistent across runs for the recurrent architecture with fixed dataset size.
Dataset scaling.
@@ 4 Results
Figure 3: Shuffled ADD: joint training fails when number structure is destroyed. Left: Learning curves for various tasks including Shuffled ADD (randomly permuted addition table). Right: Transition points—Shuffled ADD is nearly as hard as pure ADD. Unlike regular ADD, joint training with MAX+MED hurts Shuffled ADD: the mixed model reaches only $\sim$20% accuracy (third panel). This confirms that MAX+MED induces number-like embeddings incompatible with the randomized table.
To understand how language models acquire mathematical abilities, we focus on accuracy as the primary metric for observing the emergence of learning, following the approach of the emergent abilities literature (Wei et al., 2022). While some subsequent studies have questioned the concept of emergence by examining other metrics (Schaeffer et al., 2023), we argue that this critique overlooks similar patterns observed in physics: in a phase transition, only some quantities may exhibit discontinuous changes. A metric that can capture the emergence of the new phase is called the “order parameter”. Hence, we choose to use accuracy as our order parameter.
Experiments use a recurrent transformer with embedding dimension $n_{embed}$ ranging from 8 to 362, and 4 recurrence steps (we have found that using 4 recurrence steps yields the best performance). Similar to the emergence literature, we observe the total number of parameters to be the strongest indicator of accuracy, rather than $n_{embed}$ or depth. Our experimental design focuses on identifying the minimum model size sufficient for learning each task or task combination—once this threshold is exceeded, larger models reliably reach high accuracy, making further scaling unnecessary for our analysis. We primarily present results for modulo 20, with similar patterns observed for other modulos (Appendix F.5 - mod $10$ results, F.4 - mod $26$ results). Experiments with higher moduli reveal a basic scaling relationship: larger moduli require proportionally larger models and more training data to achieve comparable performance.
We find that joint training on compositional tasks substantially improves performance and reduces the parameter requirements for learning individual operations. Training on multiple arithmetic tasks helps models learn the basic building blocks needed for solving more complex problems. This shared representation not only accelerates learning but also produces abrupt transitions in performance once the necessary constituent skills are mastered (Lubana et al., 2024; Okawa et al., 2023). However, joint training is not universally beneficial, and task compatibility is more nuanced than surface-level similarity suggests. Our experiments reveal that effective transfer requires overlap in underlying computational primitives and mathematical structures, not merely similar algorithms or comparable difficulty. Pairing hard problems together often increases required model size when tasks lack shared low-level building blocks, demonstrating that joint training can actively interfere with learning rather than facilitate it. These findings extend previous work on compositional generalization (Lee et al., 2024) by systematically characterizing when multi-task learning helps or hinders algorithmic task performance, showing that compatibility depends on deep structural properties rather than intuitive task similarity.